The sum of the ages of a mother and daughter is 60 years. Also ten years ago the mother's age is 7 times the age of the of the daughter. The present age of mother is
A) 45 years
B) 55 years
C) 65 years
D) 60 years
E) None of these

Answer by ReadingBoosters(3246)

x = mom
y = daughter
x+y=60
x - 10 = 7(y-10)
Substitution
(7(y-10)+10) + y = 60
7y - 70 + 10 + y = 60
8y - 60 = 60
8y = 120
y = 15
x = 60 - 15 = 45 is mom's present age